OpenStudy (loki6871):

What is the slope-intercept form of the equation of a line that passes through (5, -4) and has a slope of 3/4 ?

OpenStudy (welshfella):

slope intercept form is y = mx + b where m = slope and b = y-intercept

OpenStudy (welshfella):

you could plug in your values into the slope point form of a straight line then rearrange to get intercept form y - y1 = m(x - x1) m = 3/4 , x1 = 5 and y1 = -4.

OpenStudy (loki6871):

thank you very much with the help